OpenStudy (anonymous):

At 7 am Joe starts sprinting at 6 miles per hour. At 7:10 am, Ken starts off after him. How fast must Ken run in order to overtake him at 7:30 am?

OpenStudy (dumbcow):

9 mph 1/3*r = 1/2*6 r = 9

OpenStudy (anonymous):

hence for joe: 6t=d (where t is time, and d is distance) for Ken: v=d/(20/60) no we know that after 30 minutes that's when ken is supposed to catch up that joe has already ran a distance of: 6*0.5=d, d=3miles substitute that back into v=d(3) then, speed of ken should be =3*3= 9 miles an hour